OpenStudy (anonymous):

30 students in a class average 80% on a certain exam. 20 others averaged 90%. what is the class average??

OpenStudy (abb0t):

Omg. You better hope you scored at least 88% on your exam cuz there is guaranteed NO CURVE.

OpenStudy (anonymous):

just answer the question plz..

OpenStudy (anonymous):

30 students got 80/100. Total score = 2400 20 students got 90/100 Total score = 1800 Total students in the class = 50 Total score of 50 students = 4200 Average = 4200/50 = 84/100 = 84%